### v1

# Dynamic Intelligence Assessment: Benchmarking LLMs on the Road to AGI with a Focus on Model Confidence

Abstract
As machine intelligence evolves, the need to test and compare the problem-solving abilities of different AI models grows. However, current benchmarks are often overly simplistic, allowing models to perform uniformly well and, making it difficult to distinguish their capabilities. Additionally, benchmarks typically rely on static question-answer pairs, which models might memorize or guess. To address these limitations, we introduce the Dynamic Intelligence Assessment (DIA), a novel methodology for testing AI models using dynamic question templates and improved metrics across multiple disciplines such as mathematics, cryptography, cybersecurity, and computer science. The accompanying DIA-Bench dataset, which includes 150 diverse and challenging task templates with mutable parameters, is presented in various formats (text, PDFs, compiled binaries, and visual puzzles). Our framework introduces four new metrics to assess a model’s reliability and confidence across multiple attempts. These metrics revealed that even simple questions are frequently answered incorrectly when posed in varying forms, highlighting significant gaps in models’ reliability. Notably, models like GPT-4o tended to overestimate their mathematical abilities, while ChatGPT-4o demonstrated better decision-making and performance through effective tool usage. We evaluated eight state-of-the-art LLMs using DIA-Bench, showing that current models struggle with complex tasks and often display unexpectedly low confidence, even with simpler questions. The DIA framework sets a new standard for assessing not only problem-solving but also a model’s adaptive intelligence and ability to assess its own limitations. The dataset is publicly available on our project’s website. https://github.com/DIA-Bench
Index Terms:
component, formatting, style, styling, insert
I Introduction
The origins of machine intelligence can be traced back to the 1950s, which simultaneously marked the need for benchmarks to evaluate its progress. The first benchmark was the famous Turing Test, introduced by Alan Turing in 1950 [1]. In 1997, IBM’s Deep Blue [2, 3] defeated Garry Kasparov in a chess match, marking a groundbreaking moment in AI, where the benchmark itself was a human expert. This victory demonstrated the potential of AI in surpassing human capabilities in complex intellectual tasks.
The introduction of neural networks and transformer-based architectures [4, 5, 6] enabled the development of powerful general-purpose models like BERT [7] and GPT-3 [8], which significantly advanced NLP capabilities. These models excelled across various tasks measured by benchmarks such as GLUE [9], SQuAD [10], and HumanEval [11]. However, these are static benchmarks, namely there is only one variant of each question and results could be memorized. Moreover, as noted by Wang et al. [12], many older benchmarks are no longer challenging enough to distinguish between newer models, as they often achieve near-perfect scores. For coding tasks, Honarvar et al. [13] explored the use of question templates to move beyond evaluating LLMs on separate, isolated problems. Apple recently extended the popular GSM8K dataset into a dynamic template-based benchmark [14]. While this allows for generating multiple variants of the same problems, the dataset remains limited to grade-school math and simple arithmetic tasks. Their findings show that LLMs struggle with even minor variations in these questions, relying more on pattern matching than true logical reasoning. In contrast, our work spans multiple disciplines and data formats, including cryptography, cybersecurity, and computer science. We move beyond accuracy-based metrics by introducing novel measures for assessing model reliability and confidence across repeated independent attempts, providing a more thorough and comprehensive evaluation of problem-solving consistency across diverse, real-world challenges.
Expanding on these, we propose Dynamic Intelligence Assessment (DIA), a testing framework that spans multiple disciplines, including cryptography, cybersecurity, and computer science. DIA moves beyond accuracy-based metrics by introducing novel measures for assessing model reliability and confidence across multiple independent instances generated from dynamic question templates. This approach allows the evaluation of LLMs consistency in problem-solving, under diverse, real-world challenges, offering a more comprehensive assessment without relying on in-context examples.
  • RQ1: What metrics can best evaluate a model’s confidence and reliability in problem-solving?
  • RQ2: How can a benchmark dataset better facilitate the accurate measurement of LLMs’ confidence and reliability in problem-solving?
  • RQ3: In problem-solving, does the use of tools impact the confidence and reliability of different models?
The main contributions can be summarized as follows:
  1. 1.
    We introduce four key metrics for assessing performance on dynamic question templates. The Reliability Score measures a model’s performance on the dataset by penalizing incorrect answers. The Task Success Rate evaluates the consistency of correct answers for a given question template. The Confidence Index captures the percentage of question templates where all $k$ versions are correctly answered, and finally, the Near Miss Score counts the number of templates where $80\%\leq$ Task Success Rate$<100\%$ for a given $k$.
  2. 2.
    We introduce the Dynamic Intelligence Assessment (DIA) framework, to provide a comprehensive assessment of AI models’ reliability and confidence. This is done by novel evaluation metrics combined with DIA-Bench, a dataset of 150 diverse, hand-crafted dynamic question templates with varying difficulty.
  3. 3.
    We tested 8 popular state-of-the-art LLMs and ranked their problem-solving reliability and confidence, providing insights on which type of tasks are particularly challenging for current models.
In summary, our work redefines how AI models are assessed for reliability, by introducing both a dynamic testing methodology with four novel metrics, shifting the focus from one-off success to consistent, reliable and confident problem-solving. The paper is organized as follows: Section III details the methodology and dataset creation, Section IV details the experimental setup and results, Section II overviews related literature, Section V discusses limitations and ethical considerations, while Section VI concludes the paper.
Figure 1: Framework for Dynamic question generation methodology, and confidence testing.
II Related Literature
II-A Metrics for Benchmarking LLMs
There is a wide range of metrics used to evaluate LLMs, each suited to different tasks. Text quality metrics such as BLEU [15], ROUGE [16], and BERTScore [17] are essential for evaluating machine translation and summarization but are less relevant to our focus. F1 score measures the average overlap between the prediction and ground truth answer. For code generation, Pass@k [11] is a key metric for the same task, whether at least one out of k generated solutions is correct.
Accuracy@k [8] measures the percentage of queries (from the entire test set) for which at least one relevant result was found within the top k results. Exact Match [10] demands an exact match with the expected output, making it particularly useful in tasks like question answering.
Robustness metrics evaluate a model’s performance under various challenging conditions [18], with Adversarial Robustness measuring the model’s resilience to attacks, such as handling unusual or malicious inputs [19]. Out-of-distribution robustness [19] measures the model’s performance on inputs significantly different from the training data, while False Refusal Rate (FRR) measures falsely rejecting benign prompts [20].
The ReCode framework [21] tests a model’s ability to maintain functionality under minor input changes. The Correctness score is the closest metric to our proposal, which measures how well an LLM answers a set of similar questions [13].
II-B Benchmark Datasets
Benchmarking LLMs has become an essential area of research across various domains, including question answering [22], code generation [23], fault localization, program repair [24], and robustness evaluation. The following is not an exhaustive list, but focuses on datasets used by industry [25], or relevant for our case.
II-B1 Question Answering Benchmarks
LLMs have been widely evaluated using question-answering (QA) benchmarks. SQuAD [10] focuses on comprehension-based QA with metrics like Exact Match (EM) and F1. CoQA [26] evaluates conversational and reading comprehension models. HotpotQA [27] tests multi-hop reasoning across multiple documents. The MINT dataset [28] benchmarks multi-round user interactions with LLMs.
CyberMetric [22] uses 10,000 multiple-choice questions to evaluate the cybersecurity knowledge of LLMs and humans. MMLU [29, 30] and MMLU-Pro [31] assess knowledge and problem-solving abilities across STEM, humanities, and social sciences. MMMU [32] benchmarks multimodal reasoning across 11.5K questions, while MMMU-Pro adds vision-only input. Lastly, DocVQA [33] focuses on visual question answering for document images.
These are just a few examples, but in general, such tests suffer from the fact, that newer LLMs are getting better in lexical knowledge and thus these datasets while useful more measuring certain domain specific knowledge, no longer allow direct comparison between LLMs, if they all simply ace the test.
II-B2 Code Generation and Problem-Solving Benchmarks
Several benchmarks evaluate LLMs’ ability to generate correct and safe code from natural language descriptions. HumanEval [11] is perhaps the most famous such dataset. MBPP [34] evaluates models on entry-level programming tasks. MATH [35] is a dataset of 12,500 challenging competition mathematics problems. MathVista [36] evaluates mathematical reasoning in visual contexts. APPS [37] uses open-access coding challenges such as Codeforces [^1] https://codeforces.com/ or Kattis [^2] https://www.kattis.com/.
II-B3 Security and Robustness Evaluation
Robustness evaluation has gained increasing importance in preventing misuse, such as jail breaking, adversarial prompting, or sensitive information leakage from training data [38, 39]. Doderlein et al. [40] developed automated operators to manipulate prompts, revealing the sensitivity of models like Copilot and Codex [11] to changes in input phrasing. Common robustness metrics include Adversarial Robustness [41] and the success rate under input perturbations, and adversarial settings.
In a small-scale study, Pearce et al. [42] demonstrated vulnerabilities in code generated by models like Copilot. The first large dataset to evaluate the security of AI-generated code was FormAI [43], Notably, it pioneered prompt templates with dynamically changing parameters to examine variations of similar programming questions. Turbulence [13] evaluates code generation with parameterized task templates but with a focus on task correctness. Wang et al. [12] introduced a framework to evolve existing datasets to be dynamic. Mirzadeh et al. [14] introduced a dynamic extension of GSM8K [44] to generate diverse variants of math problems. Their study indicated that LLMs struggle with variations in numerical values and irrelevant information, relying more on pattern matching than true logical reasoning, though the paper focuses solely on simple grade-school level arithmetic tasks. Our work expands dynamic testing by integrating diverse, dynamic, and increasingly complex challenges across multiple disciplines and data modalities.
III Methodology
First, we introduce the four evaluation metrics, followed by a discussion of the dataset creation process.
Let $t$ denote a question template with mutable parameters, where $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represents the set of such templates. For each $t\in\mathcal{T}$, the degree of freedom, denoted by $d(t)$, is the number of distinct questions that can be generated from a template. Furthermore, let $\mathcal{Q}(\mathcal{T},k)=\{q_{1},q_{2},\dots,q_{n\times k}\}$ denote the set of unique questions, where each template from $\mathcal{T}$ is used to generate $k$ different questions. In practice, we assume that $k\ll d(t)$. Let $\mathcal{S}_{\mathcal{Q}}=\{s_{1},s_{2},\dots,s_{n\times k}\}$ represent the set of solutions corresponding to $\mathcal{Q}$. To be more precise, there exists a mapping $f:\mathcal{Q}(\mathcal{T},k)\to\mathcal{S}_{\mathcal{Q}}$ such that $f(q_{i})=s_{i}$ for all $i\in\{1,2,\dots,n\times k\}$.
The Reliability Score measures the model’s performance across the dataset, with incorrect answers being heavily penalized, as shown in Formula 2. This approach is particularly useful in critical applications, where awareness of the tendency of hallucinations or incorrect responses is essential. By normalizing the score only by $k$, rather than by $k\times|\mathcal{T}|$, we can compare different instances of $\mathcal{Q}$, even when varying numbers of question instances are generated per template. Additionally, this approach still allows wrong answers to accumulate a significant negative score, providing a clear measure for human assessment about the model’s reliability over $\mathcal{T}$.
Definition 1 (Reliability Score)
The Reliability Score over a dataset $\mathcal{Q}(\mathcal{T},k)$ is calculated as:
[…]
where $\mathcal{A}_{i}$ is the score associated with answering $q_{i}$, and is defined as:
$\mathcal{A}_{i}=\begin{cases}+1&\text{if }s_{i}\text{ is returned for }q_{i},\\ 0&\text{if }q_{i}\text{ is skipped},\\ -2&\text{otherwise}.\end{cases}$ (2)
from which $[-2\times n\leq\text{RS@}(k)\leq n]$ follows.
Definition 2 (Task Success Rate)
For a given question template $t_{i}$ with $k$ generated versions, Task Success Rate measures the number of correctly solved instances, where $i\in\{1,2,\dots,n\}$.
[…]
where the value of $\mathcal{B}_{j}$ is defined as:
$\mathcal{B}_{j}=\begin{cases}+1&\text{if }s_{j}\text{ is returned for }q_{j},\\ 0&\text{if }q_{j}\text{ is skipped or answered incorrectly.}\end{cases}$ (4)
[…]
Definition 3 (Confidence Index)
represents the percentage of question templates in a dataset where, for a given template $t_{i}$, all $k$ generated queries are successfully answered,
[…]
This metric is particularly useful in critical applications, such as critical infrastructure or autonomous vehicles, where flawless task execution is essential. As $k$ increases, the metric provides stronger reassurance of the model’s reliability.
A key statistic to consider is when a model solves most queries from a question template but misses a few.
Definition 4 (Near Miss Score)
: counts the number of question templates where the model answers at least 80% but less than 100% of the instances correctly out of $k$.
[…]
When we create multiple variations of the same core problem with slight adjustments, we assess the model’s ability to consistently solve these variations. We refer to this as Local Fuzzing, where task parameters are systematically modified to evaluate the robustness of the AI model.
Definition 5 (Local Task Fuzzing)
refers to the systematic process of generating a set of $k$ instances for $\forall t\in\mathcal{T}$, by varying the mutable parameters such that $k<d(t)$.
While the four metrics were designed for evaluating dynamic datasets, they can be applied to “traditional” benchmarks with static question templates, where the same question is presented to an AI agent $k$ times. This can be particularly interesting under varying temperature settings.
DIA-Bench presents multi-modal challenges involving both visual and textual data, requiring models to process diverse formats such as PDFs and encoded files. It covers tasks including CAPTCHA solving, web security, reverse engineering, cryptography, mathematics, and logic. The multi-modal nature also stems from the fact that many questions demand specific tools like Python interpreters or Linux command-line utilities. These capabilities are not supported in LLM API calls as of today. Still, models should recognize when such tools are necessary and assess whether they can attempt a solution.
An evaluation prompt consists of a general instruction template (Figure 2), and a dynamically generated challenge description, which also always ends with “Do NOT include any explanation or additional text. ALWAYS return the answer in this XML format: ’<xml>key</xml>’”. Examples for a challenge description can be seen in Figure 34, and 5. The combination of adhering to instructions while navigating complex tasks makes DIA-Bench particularly challenging. Research consistently demonstrates that LLMs face increasing difficulties as both the number and complexity of instructions grow [14, 45, 46, 47, 48].
Figure 2: General Part of the Prompt Template. (always included)
In our dataset, the difficulty of the tasks ranges from easy to highly challenging. Although including simpler tasks, such as finding the next prime number, may seem counterintuitive, they are particularly relevant for API calls in evaluating model confidence, a topic we explore further in our discussion.
In our experiments, we tested ten Large Language Models, four of which were accessed via API calls, allowing for easy automation. We also included ChatGPT-4o, which, in some ways, acts as an orchestrator. It can run Python code, access the internet, operate within a Linux environment, and autonomously decide on the best solution. This makes it the closest candidate to general artificial intelligence as of today.
The verification module evaluates answers using an exact match, but it is flexible enough to accept minor variations, such as 5 being considered equivalent to 5.0. The evaluation process for ChatGPT-4o was time-consuming, as all queries had to be performed manually, since automating this process results in a ban from OpenAI. The 150 challenges are divided into two main categories: 50 Mathematics challenges and 100 Cybersecurity challenges. For this experiment, $k=5$ instances were generated from each of the 150 templates, resulting in a total of 750 unique questions. Figure 1 shows a visual overview of our methodology.
Figure 3: Question Template 86. A CTF-style task where ChatGPT-4o and GPT-4o API both answered ’I do not know’ five times.
Figure 4: Question Template 84: Reverse engineer. ChatGPT-4o and GPT-4o API both answered ’I do not know’ five times.
Figure 5: Question Template 76: Generate a SHA3-256 hash from the given plaintext message with adversarial prompting. GPT-4o API: Incorrect five times. ChatGPT-4o: Correct three times, Incorrect twice.
IV Discussion
TABLE I: Performance of Tested Large Language Models with $k=5$ Instances Generated for Every Task Template. (Sorted by Reliability Score).
Evaluation Metrics
Model Tool Usage Correct Skipped Wrong Reliability Score Near Miss Score Confidence Index
ChatGPT-4o Yes 418 91 241 -64 14 38.67%
GPT-4o No 112 62 576 -208 3 10.6%
Gemini-1.5-pro No 80 0 670 -252 4 6.00%
Llama 3.1 No 65 0 685 -261 1 8.6%
Gemini-1.5-flash No 60 0 690 -264 0 4.67%
qwen2.5 No 62 4 684 -261.2 0 8.2%
yi No 1 0 749 -299.4 0 0
Mistral 7B No 0 0 750 -300 0 0
We evaluated eight models: GPT-4o, Gemini-1.5-pro, Gemini-1.5-flash, Llama 3.1, qwen2.5, Mistral 7B, and yi, along with one chat model equipped with tool-using capabilities, ChatGPT-4o. Notably, existing literature often overlooks whether the LLMs under examination possess tool-usage capabilities, and many benchmarks fail to specify whether the chat or API version of a model is employed.
In this section, we focus primarily on comparing the OpenAI models. First, because they are the top performers, and second, one can use tools while the other cannot, offering the most compelling case for comparison. While the initial assumption might be that ChatGPT-4o performs better solely because of its tool-using capabilities, a more nuanced analysis suggests that the gap between the two models is not purely a function of tool availability.
IV-A An Enthusiasm for Mathematics
One of the most striking observations is GPT-4o API’s eagerness to engage with mathematical tasks, even though it consistently fails to solve them correctly, as shown in Figure 6. This behaviour could stem from the fact that GPT-4o API and ChatGPT-4o share the same underlying architecture. In mathematics, GPT-4o consistently assumes that it has the capacity to solve the problems. This aligns with the observation made by Mirzadeh et al. [14], where large language models (LLMs) often rely on sophisticated pattern matching rather than true logical reasoning.
GPT-4o API’s inability to skip tasks reflects a broader issue in its self-assessment. Despite prompts encouraging it to skip problems it cannot solve, GPT-4o API rarely does so. This indicates a lack of meta-cognitive awareness in gauging its limitations, particularly when tools are required to solve a task. It continues to attempt math problems, even though it lacks the necessary tool-using capabilities to succeed.
Figure 6: Comparison of (a) GPT-4o results and (b) ChatGPT-4o results in Cybersecurity (100 templates, 500 questions generated) and Mathematics with (50 templates, 250 questions generated)
IV-B Cybersecurity vs. Mathematics: Different Approaches in Skipping
The contrast between the models’ behavior on mathematical tasks versus cybersecurity tasks provides additional insights. GPT-4o API, which eagerly attempts mathematical problems, is far more reserved in cybersecurity tasks, skipping questions almost 20 times more frequently (3 vs. 59 as shown in Figure 6). In cybersecurity tasks, both models tend to skip more frequently, suggesting that these tasks invoke a higher degree of self-reflection and caution. This behaviour indicates that both models are capable of critical evaluation, particularly when the task is complex enough to warrant skipping. Overall, ChatGPT-4o demonstrates an improved ability to assess when a task exceeds its capabilities, as evidenced by its higher tendency to skip difficult tasks.
IV-C Limitations in Current Metrics
In the traditional setting, skipping is an inferior approach to hallucination or guessing. To underline this, we should examine ChatGPT-4o’s performance using a traditional metric like Pass@k. On the DIA-Bench, ChatGPT-4o achieved an impressive 73.3% Pass@5, meaning it was able to provide a correct answer to $73.3\%$ of the question templates when given five attempts. At the same time it was only able to ace all five instances for $38.67\%$ of the templates. The same goes for GPT-4o API, with Pass@5=21.3% and Conf@5=10.6%. The significant gap between the Pass@k score and the Confidence Index highlights the importance of reliability metrics.
Moreover, if we would rely on exact match for $k=1$, as is common with static datasets, at least a third of the tasks on DIA-Bench would have a minimum of 1 in 5 chance of being answered correctly through trial and error. This artificially boosts the success rate but does not accurately reflect a model’s reliability and confidence.
IV-D A lack of Confidence
ChatGPT-4o recorded a high near-miss score, achieving 4 out of 5 on 14 templates. We anticipate that as $k$ increases, the near-miss score would rise, lowering the confidence index. During early experimentation with prompt templates, we observed that models often approached tasks with varying strategies. For example, when tasked with adding two smaller numbers, ChatGPT-4o might not immediately write and execute Python code—considering the task trivial—but instead produce an answer close to the correct one. For more complex tasks, it demonstrated this behaviour less.
ChatGPT-4o completely skipped 9 out of 150 templates, but inconsistently skipped at least one instance in 32 other templates, revealing gaps in its decision-making. This suggests that its self-assessment mechanism isn’t fully reliable, as it sometimes engages with tasks it later recognizes as too difficult. Despite this, ChatGPT-4o outperforms its API counterparts, which is expected since API models lack tools and code execution capabilities. Interestingly, both GPT-4o and ChatGPT-4o exhibit similar “taste” when attempting task categories, with GPT-4o taking on math problems enthusiastically despite lacking tools. In cybersecurity tasks, GPT-4o and ChatGPT-4o display a strong correlation in their skipping patterns—when GPT-4o skips a template, ChatGPT-4o often does the same for at least some instances of that template.
Other API models rarely utilized the skip option and attempted to answer every question. This pattern, evident in Table I, underscores that most models are primarily engaged in pattern matching and remain distant from achieving artificial general intelligence (AGI). The results reveal a substantial gap in problem-solving performance between models with tool-using capabilities and those without. LLMs without tools tend to hallucinate more often, failing to complete tasks accurately, which is reflected in their lower reliability scores.
V Limitations and Threats to Validity
The accuracy and reliability of benchmarking outcomes depend on the quality of the generator scripts, and despite thorough proof checking on our part, some errors or edge cases may persist. If any issues are found, please report them by opening a GitHub issue.
Certain task templates might not be as well represented as others, leading to an incomplete evaluation of the LLM’s overall performance. In our study, we focused on computer science and mathematics, omitting other disciplines to concentrate our resources. We aim to expand this, and the used data modalities.
We did not utilize o1-preview or other chat models due to their high time and cost requirements, which would have impacted the feasibility of this work.
VI Conclusion
In this study, we introduce the Dynamic Intelligence Assessment (DIA) framework along with the DIA-Bench dataset to test the problem-solving reliability and confidence of large language models (LLMs) for mathematics and computer science questions. We also identified four key metrics for evaluating LLMs reliability and confidence. Our work addressed three key research questions:
  • RQ1: What metrics can best evaluate a model’s confidence and reliability in problem-solving? Answer: Traditional metrics often provide a misleading picture of a model’s true performance, as they don’t account for consistency. We introduce four metrics: the Reliability Score, which penalizes incorrect answers, the Task Success Rate (TSR) for measuring consistency across repeated instances, the Confidence Index to assess performance across variations of the same task, and the Near Miss Score for tasks where the model almost but not fully succeeds.
  • RQ2: How can a benchmark dataset better facilitate the accurate measurement of LLMs’ confidence and reliability in problem-solving?
    Answer: By using hand-crafted dynamic question templates, hosting simpler tasks to very difficult ones across various formats (text, PDFs, compiled binaries, etc.). This, combined with the use of the proposed evaluation metrics, offers a more rigorous comparison of models’ abilities compared to static, non-multimodal, or overly easy benchmarks.
  • RQ3: In problem-solving, does the use of tools impact the confidence and reliability of different models?
    Answer: Yes, the use of tools significantly enhances both the confidence and reliability of models in problem-solving. Tool-using models like ChatGPT-4o not only perform better on complex tasks requiring precise computations, but they also demonstrate more consistent and reliable results across various task types. Interestingly, there is a strong correlation between the GPT models in their tendency to attempt to solve a task, where non-tool users are more prone to errors and hallucinations.
Our findings demonstrate that while current LLMs have made significant progress, there remain challenges in achieving consistent and reliable problem-solving, particularly for models without tool-using capabilities. DIA-Bench provides a valuable resource for further improving LLM evaluation and development towards more adaptable and reliable AI systems.